OpenStudy (anonymous):

Find a so that f(x) {7x+10 if x < or = 3 {-2x+a if x>3 is continuous at x=3.

OpenStudy (anonymous):

Find a so that f(x) {7x+10 if x \[\le\] 3 {-2x+a if x>3 is continuous at x=3.

OpenStudy (nikvist):

\[f(3)=7\cdot 3+10=31\]\[f(3)=\lim_{x\rightarrow 3}(-2x+a)=-6+a=31\]\[a=37\]